# Imperial College London

### v1

@@ Abstract
Analog computing has been recently revived due to its potential for energy-efficient and highly parallel computations. In this paper, we investigate analog computers that linearly process microwave signals, named microwave linear analog computers (MiLACs), and their applications in signal processing for communications. We model a MiLAC as a multiport microwave network with tunable impedance components, which enables the execution of mathematical operations by reconfiguring the microwave network and applying input signals at its ports. We demonstrate that a MiLAC can efficiently compute the linear minimum mean square error (LMMSE) estimator, widely used in multiple-input multiple-output (MIMO) communications beamforming and detection, with remarkably low computational complexity, unachievable through digital computing. Specifically, the LMMSE estimator can be computed with complexity growing with the square of its input size, rather than the cube, with revolutionary applications to gigantic MIMO beamforming and detection.
@@
Analog computing, linear minimum mean square error (LMMSE) estimator, microwave networks.
@@ I Introduction
Computing devices play a pivotal role in our society, serving as the backbone of the technological advancement. Modern computers involve digital electronic circuits operating on discrete values (binary) to encode and manipulate data. Digital circuits and signals are widespread today because of their numerous benefits over their analog counterparts [1], such as versatility, more precision and noise robustness, flexibility in storing and processing data, and security.
@@
Analog computers based on optical systems, or optical analog computers, leverage the speed and parallelism of light to process data, represented as optical signals, at an exceptionally high rate [3, 4]. Tasks related to image processing and machine learning inference have been performed through optical analog computers, as reviewed in [9]. Although optical analog computers have demonstrated significant advantages in low-power and ultra-fast processing, they typically rely on large-scale architectures, as their size cannot be reduced below the operational wavelength.
@@
To realize miniaturized analog computers at scales smaller than the operational wavelength, metamaterials have emerged as artificial materials designed to have unusual electromagnetic (EM) properties [5, 6]. Analog computers based on metamaterials have demonstrated the possibility of solving mathematical equations [10], including multiple equations in parallel exploiting input signals at different frequencies [11]. Particular interest has been dedicated to the 2-dimensional counterpart of metamaterials, namely metasurfaces, further reducing the dimensions of the computing devices. Metasurfaces can be engineered to apply specific transfer functions to the spatial and frequency components of the incident wave, including differentiation, integration, and convolution [5, 6]. Metasurfaces are also expected to play a crucial role in future wireless communications due to their ability to perform computations directly in the wave domain [12, 13, 14].
@@
Besides metamaterials and metasurfaces, resistive memory arrays have also emerged as a prominent technology for building modern analog computing devices [7, 8]. Particularly noteworthy are crosspoint memory arrays, in which resistive memory cells are located at the intersections of horizontal and vertical lines in a matrix shape. The values stored in these memory cells can naturally represent the entries of a matrix, enabling the memory array to perform matrix operations, such as matrix-vector multiplications, directly in the analog domain [15, 16]. The parallelism offered by resistive memory arrays in computing matrix-vector multiplications has been used to advance future multiple-input multiple-output (MIMO) communications in [17, 18].
@@
Several communities have developed modern analog computing solutions to accelerate signal processing operations [3]-[11], [15, 16] as well as to foster wireless communications [12, 13, 14, 17, 18]. However, the efforts of these communities are often siloed, with limited inter-disciplinarity, and this fragmentation has led to a lack of a unified model and shared theoretical understanding. In this paper, we aim to fill this gap by focusing on analog computers that linearly process microwave signals. We develop the model of these analog computers by using rigorous microwave theory, characterize their input-output relationship, and explore their application to signal processing for wireless communications.
@@
The contributions of this paper can be summarized as follows. First, we derive a general model for an analog computer that processes microwave signals linearly, i.e., satisfying the superposition principle, which we denote as microwave linear analog computer (MiLAC). A MiLAC can be regarded as a microwave network having multiple ports, where input signals are applied and output signals are read. This is the first effort in providing a theoretical framework unifying previous works [3]-[18]. Second, we characterize the output of a MiLAC as a function of the input signals and the impedance components constituting its microwave network. The obtained expressions involve matrix-matrix products and matrix inversions, which are computationally expensive operations in the digital domain, as their complexity scales cubically with the matrix size. Nevertheless, they can be instantly computed by a MiLAC once the microwave network is properly reconfigured and the input signals are applied. Third, we show that a MiLAC can be used as a computing device to efficiently calculate the linear minimum mean square error (LMMSE) estimator, widely adopted in wireless communications. Compared to a digital computer, a MiLAC can calculate the LMMSE estimator with a reduction in computational complexity by a factor of $2.5\times 10^{4}$.
@@
Notation: Vectors and matrices are denoted with bold lower and bold upper letters, respectively. Scalars are represented with letters not in bold font. $\mathbf{a}^{T}$, $\mathbf{a}^{H}$, $[\mathbf{a}]_{i}$, and $\|\mathbf{a}\|_{2}$ refer to the transpose, conjugate transpose, $i$th element, and $l_{2}$-norm of a vector $\mathbf{a}$, respectively. $\mathbf{A}^{T}$, $\mathbf{A}^{H}$, and $[\mathbf{A}]_{i,k}$ refer to the transpose, conjugate transpose, and $(i,k)$th element of a matrix $\mathbf{A}$, respectively. $\mathbb{C}$ denotes the complex number set. $j=\sqrt{-1}$ denotes the imaginary unit. $\mathbf{I}_{N}$ and $\mathbf{0}_{N}$ denote the identity matrix and the all-zero matrix with dimensions $N\times N$, respectively, and $\mathbf{0}_{M\times N}$ denote the all-zero matrix with dimensions $M\times N$.
@@ II-A Microwave Network
The circuit implementing a MiLAC is in general a linear microwave network, following the definition of MiLAC. According to microwave network theory [19, Chapter 4], a linear microwave network having a certain number of ports $P$ can be characterized by its admittance matrix $\mathbf{Y}\in\mathbb{C}^{P\times P}$, which linearly relates the voltages and currents at the $P$ ports through
@@
where $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}\in\mathbb{C}^{P\times 1}$ and $\mathbf{i}=[i_{1},\ldots,i_{P}]^{T}\in\mathbb{C}^{P\times 1}$ are the voltage and current vectors at the ports, with $v_{p}\in\mathbb{C}$ and $i_{p}\in\mathbb{C}$ being the voltage and current at the $p$th port, for $p=1,\ldots,P$.
@@ II-B Input Signals
Assume that the MiLAC receives the input on the first $N$ ports of its reconfigurable microwave network, with $N<P$. Note that this comes with no loss of generality, as the ports can always be reordered to ensure the input is applied on the first $N$ ports. In addition, if an input needs to be applied on an internal node of the microwave network, an additional port can be opened, making this model general. The input is applied by $N$ voltage sources with their series impedance $Z_{0}$, e.g., $Z_{0}=50~{}\Omega$, or, equivalently, series admittance $Y_{0}=Z_{0}^{-1}$, as shown in Fig. 1. Denoting as $u_{n}\in\mathbb{C}$ the voltage of the $n$th voltage source, for $n=1,\ldots,N$, the input vector is given by $\mathbf{u}=[u_{1},\ldots,u_{N}]^{T}\in\mathbb{C}^{N\times 1}$. Following Ohm’s law, the input voltage $u_{n}$ is related to the voltage $v_{n}$ and current $i_{n}$ at the $n$th port by $i_{n}=Y_{0}(u_{n}-v_{n})$, for $n=1,\ldots,N$, giving
@@ II-C Output Signals
The output of the MiLAC is read on all the $P$ ports of its reconfigurable microwave network, i.e., is given by the vector $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}$. If an output needs to be read on an internal node of the microwave network, an additional port can be added, confirming the generality of our model. Since the number of input ports $N$ is assumed to be lower than the number of total ports $P$, we can write $P=N+M$, where $M$ is the number of ports with no input signal. To read the output voltage on these $M$ ports without causing undesired reflections within the microwave network, these $M$ ports must be perfectly matched to $Z_{0}$, i.e., they are terminated to ground via an impedance $Z_{0}$, or, equivalently, an admittance $Y_{0}=Z_{0}^{-1}$, as shown in Fig. 1. Thus, according to Ohm’s law, the voltage and current at the $(N+m)$th port are related by $i_{N+m}=-Y_{0}v_{N+m}$, for $m=1,\ldots,M$, where the negative sign arises from the direction of the current $i_{N+m}$. This yields
@@ II-D Tunable Admittance Components
The microwave network of the MiLAC must be implemented with tunable admittance components, such that its admittance matrix $\mathbf{Y}$ is reconfigurable, allowing the computation of different operations in the analog domain. To realize a MiLAC with maximum flexibility, i.e., whose admittance matrix $\mathbf{Y}$ can be reconfigured to any complex matrix with dimension $P\times P$, we can implement its microwave network with $P^{2}$ tunable admittance components connecting each port to ground and to all other ports. Specifically, port $k$ is connected to ground through an admittance $Y_{k,k}\in\mathbb{C}$, for $k=1,\ldots,P$, and port $k$ is connected to port $i$ through an admittance $Y_{i,k}\in\mathbb{C}$, $\forall i\neq k$. An example of such a MiLAC with maximum flexibility is reported in Fig. 2, where we have $P=4$ ports and $N=2$ input signals.
This topology of tunable admittance components allows us to arbitrarily reconfigure the admittance matrix $\mathbf{Y}$ since its $(i,k)$th element is given by
@@
for $i,k=1,\ldots,P$, following microwave network theory [19, Chapter 4]. By inverting (4), we can derive the expression of the tunable admittance components as a function of any given admittance matrix $\mathbf{Y}$ as
@@
for $i,k=1,\ldots,P$, showing that the $P^{2}$ tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ can be adjusted to allow $\mathbf{Y}$ to assume any arbitrary value.
Figure 2: Representation of a $4$-port MiLAC with input on $2$ ports.

### v2

# Analog Computing for Signal Processing and Communications – Part I: Computing with Microwave Networks

@@ Abstract
Analog computing has been recently revived due to its potential for energy-efficient and highly parallel computations. In this two-part paper, we explore analog computers that linearly process microwave signals, named microwave linear analog computers (MiLACs), and their applications in signal processing and communications. In Part I of this paper, we model a MiLAC as a multiport microwave network with tunable impedance components, enabling the execution of mathematical operations by reconfiguring the microwave network and applying input signals at its ports. We demonstrate that a MiLAC can efficiently compute the linear minimum mean square error (LMMSE) estimator and matrix inversion, with remarkably low computational complexity. Specifically, a matrix can be inverted with complexity growing with the square of its size. We also show how a MiLAC can be used jointly with digital operations to implement sophisticated algorithms such as the Kalman filter. To enhance practicability, we propose a design of MiLAC based on lossless impedance components, reducing power consumption and eliminating the need for costly active components. In Part II of this paper, we investigate the applications of MiLACs in wireless communications, highlighting their potential to enable future wireless systems by executing computations and beamforming in the analog domain.
@@
Analog computing, Kalman filter, linear minimum mean square error (LMMSE) estimator, matrix inversion, microwave networks.
@@ I Introduction
Computing devices play a pivotal role in our society, serving as the backbone of the technological advancement. As our world becomes increasingly technology-driven and interconnected, we continuously use computing devices to process data, automate tasks, solve problems, and exchange information. Modern computers involve digital electronic circuits operating on discrete values (binary) to encode and manipulate data. Digital circuits and signals are widespread today because of their numerous benefits over their analog counterparts [2], such as versatility, more precision and noise robustness, flexibility in storing and processing data, and security.
@@
Analog computers based on optical systems, or optical analog computers, leverage the speed and parallelism of light to process data, represented as optical signals, at an exceptionally high rate [4, 5]. Optical analog computers have been designed to efficiently perform linear operations [10] as well as compute the inverse of a given matrix [11]. More complicated tasks have also been performed through optical analog computers, specifically related to image processing and machine learning inference, as reviewed in [12]. In [13], a 2-dimensional spatial differentiator has been built to transform the input optical signal into its second-order derivative, enabling edge detection within an image. Machine learning inference in the analog domain has been demonstrated in [14], where the functioning of a neural network (NN) has been emulated via the propagation of an optical signal through multiple diffractive layers, each acting as a layer of a NN. Diffractive optics has also been used to build a hybrid optical-electronic convolutional NN, in which the convolution operation is performed in the analog domain through a phase mask [15]. Although optical analog computers have demonstrated significant advantages in low-power and ultra-fast processing, they typically rely on large-scale architectures, as their size cannot be reduced below the operational wavelength.
@@
To realize miniaturized analog computers at scales smaller than the operational wavelength, metamaterials have emerged as artificial materials designed to have unusual electromagnetic (EM) properties [6, 7]. Because of their artificially induced properties, metamaterials allow us to manipulate EM signals within sub-wavelength scales. Analog computers based on metamaterials have demonstrated the possibility of solving mathematical equations [16], including multiple equations in parallel exploiting input signals at different frequencies [17]. Particular interest has been dedicated to the 2-dimensional counterpart of metamaterials, namely metasurfaces, further reducing the dimensions of the computing devices. Metasurfaces can be engineered to apply specific transfer functions to the spatial and frequency components of the incident wave, including differentiation, integration, and convolution [6, 7]. To perform these mathematical operations, various metasurfaces technologies have been developed, such as space-time digital metasurfaces [18] and nonlocal metasurfaces having an angular-dependent response [19, 20]. Beyond linear transformations [21], metasurfaces have been designed to perform edge detection and pattern recognition, which are crucial operations in image processing [22, 23]. The capability of metasurfaces to perform matrix-vector multiplications in the analog domain has been exploited to implement programmable NNs performing recognition tasks [24] and reinforcement learning tasks [25]. Metasurfaces are also expected to play a crucial role in future wireless communications due to their ability to perform computations directly in the wave domain [26]. Possible applications of computing metasurfaces in wireless communications include performing sensing tasks and enhancing physical layer security [27, 28].
@@
Besides metamaterials and metasurfaces, resistive memory arrays have also emerged as a prominent technology for building modern analog computing devices [8, 9]. Particularly noteworthy are crosspoint memory arrays, in which resistive memory cells are located at the intersections of horizontal and vertical lines in a matrix shape. The values stored in these memory cells can naturally represent the entries of a matrix, enabling the memory array to perform matrix operations, such as matrix-vector multiplications, directly in the analog domain [29, 30]. This approach, referred to as in-memory computing or analog matrix computing, has demonstrated the capability to execute matrix-vector multiplications with arbitrarily high precision while consuming less energy compared to digital computing [31]. The parallelism offered by resistive memory arrays in computing matrix-vector multiplications has been harnessed to effectively implement and accelerate NNs [32, 33, 34]. Furthermore, analog matrix computing via resistive memory arrays has been used to advance future multiple-input multiple-output (MIMO) communications [35, 36].
@@
Several communities have developed modern analog computing solutions to accelerate signal processing operations [4]-[25], [29]-[34] as well as to foster wireless communications [26, 27, 28, 35, 36]. However, the efforts of these communities are often siloed, with limited inter-disciplinarity, and this fragmentation has led to a lack of a unified model and shared theoretical understanding. In this two-part paper, we aim to fill this gap, by focusing on analog computers that linearly process microwave signals. We develop the model of these analog computers by using rigorous microwave theory, characterize their input-output relationship, and explore their application to signal processing (in Part I) and wireless communications (in Part II). Specifically, the contributions of Part I of this paper are as follows.
@@
First, we derive a general model for an analog computer that processes microwave signals satisfying the superposition principle, which we denote as microwave linear analog computer (MiLAC). A MiLAC can be regarded as a microwave network having multiple ports, where input signals are applied and output signals are read. Once the input signals are applied, they propagate at light speed within the microwave network generating the output signals. In this way, specific operations can be instantly computed by the MiLAC, requiring no analog-digital conversion or digital computations. This is the first effort in providing a theoretical framework unifying previous works [4]-[36].
Second, we characterize the output of a MiLAC as a function of the input signals and the impedance components constituting its microwave network. Specifically, we derive the expression of the output signals by using microwave theory, and regarding the MiLAC as a multiport microwave network. The obtained expressions involve matrix-matrix products and matrix inversions, which are computationally expensive operations in the digital domain, as their complexity scales cubically with the matrix size. Nevertheless, they can be instantly computed by a MiLAC once the microwave network is properly reconfigured and the input signals are applied. While it is known that specific matrix operations can be computed in the analog domain [4]-[36], we provide the general expression for the functions computable with a MiLAC.
Third, we show that a MiLAC can be used as a computing device to efficiently calculate popular mathematical operations, namely the linear minimum mean square error (LMMSE) estimator and the inverse of a given matrix. Compared to a digital computer, a MiLAC can calculate these two operations with a reduction in computational complexity by a factor of $2.5\times 10^{4}$ and $5.5\times 10^{3}$, respectively. We define a set of primitive operations computable with a MiLAC, and show how they can be sequentially used jointly with digital operations to efficiently implement more sophisticated algorithms, such as the Kalman filter. Specifically, a Kalman filter can be implemented with a MiLAC by reducing the computational complexity by $1.1\times 10^{4}$ times. The sequential use of primitive operations computed in the analog domain is less explored in literature [4]-[36], and can further extend the set of possible functions implementable with analog computers.
Fourth, we propose a practical design to implement a MiLAC, relying on a microwave network solely constituted of lossless tunable impedance components, i.e., with purely imaginary values instead of complex. The use of such components is expected to reduce losses and avoid costly active radio-frequency (RF) components, providing the initial guidelines for the practical implementation of a MiLAC. We show that such a MiLAC made of imaginary impedance components can execute the same operations as a MiLAC built with complex impedance components.
Organization: In Section II, we present the general model a MiLAC. In Section III, we analyze a MiLAC and characterize its output as a function of its input. In Section IV, we show how a MiLAC can efficiently compute the LMMSE estimator and invert a matrix. In Section V, we show that a MiLAC can also efficiently compute the Kalman filter. In Section VI, we propose a practical implementation of a MiLAC, based on lossless and reciprocal tunable impedance components. Finally, Section VII concludes Part I of this paper.
@@
Notation: Vectors and matrices are denoted with bold lower and bold upper letters, respectively. Scalars are represented with letters not in bold font. $\Re\{a\}$, $\Im\{a\}$, and $|a|$ refer to the real part, imaginary part, and absolute value of a complex scalar $a$, respectively. $\mathbf{a}^{T}$, $\mathbf{a}^{H}$, $[\mathbf{a}]_{i}$, and $\|\mathbf{a}\|_{2}$ refer to the transpose, conjugate transpose, $i$th element, and $l_{2}$-norm of a vector $\mathbf{a}$, respectively. $\mathbf{A}^{T}$, $\mathbf{A}^{H}$, $[\mathbf{A}]_{i,k}$, $[\mathbf{A}]_{:,k}$, and $\|\mathbf{A}\|_{F}$ refer to the transpose, conjugate transpose, $(i,k)$th element, $k$th column, and Frobenius norm of a matrix $\mathbf{A}$, respectively. $\mathbb{R}$ and $\mathbb{C}$ denote the real and complex number sets, respectively. $j=\sqrt{-1}$ denotes the imaginary unit. $\mathbf{I}_{N}$ and $\mathbf{0}_{N}$ denote the identity matrix and the all-zero matrix with dimensions $N\times N$, respectively, and $\mathbf{0}_{M\times N}$ denote the all-zero matrix with dimensions $M\times N$.
@@ II-A Microwave Network
The circuit implementing a MiLAC is in general a linear microwave network, following the definition of MiLAC. According to microwave network theory [37, Chapter 4], a linear microwave network having a certain number of ports $P$ can be characterized by its impedance matrix $\mathbf{Z}\in\mathbb{C}^{P\times P}$, which linearly relates the voltages and currents at the $P$ ports through
$\mathbf{v}=\mathbf{Z}\mathbf{i},$ (1)
where $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}\in\mathbb{C}^{P\times 1}$ and $\mathbf{i}=[i_{1},\ldots,i_{P}]^{T}\in\mathbb{C}^{P\times 1}$ are the voltage and current vectors at the ports, with $v_{p}\in\mathbb{C}$ and $i_{p}\in\mathbb{C}$ being the voltage and current at the $p$th port, for $p=1,\ldots,P$. Alternatively, a linear microwave network with $P$ ports can also be represented by its admittance matrix $\mathbf{Y}\in\mathbb{C}^{P\times P}$, given by $\mathbf{Y}=\mathbf{Z}^{-1}$, relating the currents and voltages at the $P$ ports as
@@
equivalently to (1) [37, Chapter 4]. Note that impedance and admittance matrices are equivalent representations of a linear microwave network, allowing its analysis regardless of its internal circuit topology.
@@ II-B Input Signals
Assume that the MiLAC receives the input on the first $N$ ports of its reconfigurable microwave network, with $N\leq P$. Note that this comes with no loss of generality, as the ports can always be reordered to ensure the input is applied on the first $N$ ports. In addition, if an input needs to be applied on an internal node of the microwave network, an additional port can be opened, making this model general. The input is applied by $N$ voltage sources with their series impedance $Z_{0}$, e.g., $Z_{0}=50~{}\Omega$, or, equivalently, series admittance $Y_{0}=Z_{0}^{-1}$, as shown in Fig. 1. Denoting as $u_{n}\in\mathbb{C}$ the voltage of the $n$th voltage source, for $n=1,\ldots,N$, the input vector is given by $\mathbf{u}=[u_{1},\ldots,u_{N}]^{T}\in\mathbb{C}^{N\times 1}$. Following Ohm’s law, the input voltage $u_{n}$ is related to the voltage $v_{n}$ and current $i_{n}$ at the $n$th port by $i_{n}=Y_{0}(u_{n}-v_{n})$, for $n=1,\ldots,N$, giving
@@ II-C Output Signals
The output of the MiLAC is read on all the $P$ ports of its reconfigurable microwave network, i.e., is given by the vector $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}$. If an output needs to be read on an internal node of the microwave network, an additional port can be added, confirming the generality of our model. Since the number of input ports $N$ is never higher than the number of total ports $P$, we can write $P=N+M$, where $M$ is the number of ports with no input signal. To read the output voltage on these $M$ ports without causing undesired reflections within the microwave network, these $M$ ports must be perfectly matched to $Z_{0}$, i.e., they are terminated to ground via an impedance $Z_{0}$, or, equivalently, an admittance $Y_{0}=Z_{0}^{-1}$, as shown in Fig. 1. Thus, according to Ohm’s law, the voltage and current at the $(N+m)$th port are related by $i_{N+m}=-Y_{0}v_{N+m}$, for $m=1,\ldots,M$, where the negative sign arises from the direction of the current $i_{N+m}$. This yields
@@ II-D Tunable Admittance Components
The microwave network of the MiLAC must be implemented with tunable admittance components, such that its impedance matrix $\mathbf{Z}$ and admittance matrix $\mathbf{Y}$ are reconfigurable, allowing the computation of different operations in the analog domain. As impedance and admittance matrices are equivalent representations of a microwave network, we use the admittance matrix $\mathbf{Y}$ representation in this study since it can be directly related to the tunable components of the microwave network, as will become clear in the remainder of this section.
Figure 2: Representation of a $4$-port MiLAC with input on $2$ ports.
To realize a MiLAC with maximum flexibility, i.e., whose admittance matrix $\mathbf{Y}$ can be reconfigured to any complex matrix with dimension $P\times P$, we can implement its microwave network with $P^{2}$ tunable admittance components connecting each port to ground and to all other ports. Specifically, port $k$ is connected to ground through an admittance $Y_{k,k}\in\mathbb{C}$, for $k=1,\ldots,P$, and port $k$ is connected to port $i$ through an admittance $Y_{i,k}\in\mathbb{C}$, $\forall i\neq k$. An example of such a MiLAC with maximum flexibility is reported in Fig. 2, where we have $P=4$ ports and $N=2$ input signals.
This topology of tunable admittance components allows us to arbitrarily reconfigure the admittance matrix $\mathbf{Y}$, as it can be shown by following the definition of $\mathbf{Y}$ in (2). According to (2), $[\mathbf{Y}]_{i,k}$ is given by applying a voltage $v_{k}$ to port $k$, short-circuiting all the other ports, i.e., $v_{p}=0$, $\forall p\neq k$, measuring the current $i_{i}$ entering at port $i$, and computing the ratio
$\left[\mathbf{Y}\right]_{i,k}=\left.\frac{i_{i}}{v_{k}}\right|_{v_{p}=0,\forall p\neq k},$ (5)
for $i,k=1,\ldots,P$ [37, Chapter 4]. Following (5), the entries $\left[\mathbf{Y}\right]_{i,k}$ can be computed as a function of the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ by separately studying the two cases of off-diagonal entries $\left[\mathbf{Y}\right]_{i,k}$ with $i\neq k$ and diagonal entries $\left[\mathbf{Y}\right]_{i,k}$ with $i=k$. First, when $i\neq k$, the ratio in (5) can be determined by studying the circuit in Fig. 3(a), since all the other admittance components in the microwave network do not impact the current $i_{i}$. Thus, we obtain the off-diagonal entry $\left[\mathbf{Y}\right]_{i,k}$ as $\left[\mathbf{Y}\right]_{i,k}=-Y_{i,k}$ following Ohm’s law, where the negative sign arises from the direction of the current $i_{i}$. Second, when $i=k$, the ratio in (5) can be determined by studying the circuit in Fig. 3(b), since all the other admittance components do not influence $i_{k}$. In this case, the diagonal entry $\left[\mathbf{Y}\right]_{k,k}$ is given by the parallel of the $P$ admittance components $Y_{1,k},\ldots,Y_{P,k}$, i.e., $\left[\mathbf{Y}\right]_{k,k}=\sum_{p=1}^{P}Y_{p,k}$. By summarizing the two examined cases, we can write
@@
for $i,k=1,\ldots,P$, giving the expression of each entry of the admittance matrix $\mathbf{Y}$ as a function of the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$. By inverting (6), we can derive the expression of the tunable admittance components as a function of any given admittance matrix $\mathbf{Y}$ as
@@
for $i,k=1,\ldots,P$, showing that the $P^{2}$ tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ can be adjusted to allow $\mathbf{Y}$ to assume any arbitrary value [^1] Equation (7) clarifies the direct relationship between the entries of the admittance matrix $\mathbf{Y}$ and the tunable components of the microwave network. Note that closed-form expressions are not available to relate the entries of the impedance matrix $\mathbf{Z}$ and the tunable components..
Note that it is also possible to implement a MiLAC with a reduced number of tunable admittance components. However, such a MiLAC will experience reduced flexibility since the microwave network cannot be reconfigured to take any arbitrary admittance matrix $\mathbf{Y}$, as a consequence of the reduced circuit complexity. In this work, we consider a MiLAC having maximum flexibility and leave the analysis of the trade-off between circuit complexity and flexibility as future research.
Figure 3: Circuit to be analyzed to compute (a) the off-diagonal entry $[\mathbf{Y}]_{i,k}$ and (b) the diagonal entry $[\mathbf{Y}]_{k,k}$ as a function of the tunable admittance components.